OpenStudy (anonymous):

whats the standard form of a line passing through the points (2,-3) and (4,2)

OpenStudy (texaschic101):

you need to first find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (2,-3) x1 = 2 and y1 = -3 (4,2) x2 = 4 and y2 = 2 now we sub slope(m) = (2 - (-3) / (4 - 2) can you finish and find the slope ?

OpenStudy (texaschic101):

let me know when your ready to participate :)

OpenStudy (amistre64):

@texaschic101 gonna see if you recall what standard form looks like :)

OpenStudy (texaschic101):

Ax + By = C is standard form

OpenStudy (anonymous):

hold on

OpenStudy (amistre64):

very good ;)

OpenStudy (texaschic101):

you figure the slope out yet ??

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

slope(m) = (2 - (-3) / (4 - 2) slope(m) = (2 + 3) / 2 slope(m) = 5/2 you understand so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

no we will use y = mx + b form.... slope(m) = 5/2 you can use either of your points...I will use (4,2) x = 4 and y = 2 now we sub and solve for b y = mx + b 2 = 5/2(4) + b 2 = 10 + b 2 - 10 = b -8 = b so our equation is : y = 5/2x - 8 still with me or do you have any questions ?

OpenStudy (texaschic101):

now we need to turn this into Ax + By = C y = 5/2x - 8 -- multiply the entire equation by 2 2y = 5x - 16 -- subtract 5x from both sides -5x + 2y = -16 -- multiply by -1 to make x positive 5x - 2y = 16 ==> standard form